# Efficient Ensemble Conditional Independence Test Framework for Causal Discovery

### v1

@@ 4 Experiments
In this section, we comprehensively evaluate the effectiveness of E-CIT. We first demonstrate its ability to reduce computational costs while maintaining performance (Section 4.1), followed by its broad applicability across different CIT methods (Section 4.2) and strong performance on real-world datasets (Section 4.3). Additionally, we apply E-CIT in Causal Discovery (Section 4.4). Other results, including the impact of subset size and the advantages of our p-value combination method for CIT, can be found in Appendix E.6 and E.7.
We conduct our synthetic experiments under the post-nonlinear model, following the setup of prior works (Zhang et al., 2011; Doran et al., 2014; Bellot & van der Schaar, 2019; Scetbon et al., 2022; Li et al., 2023a; b). Specifically, we consider the null hypothesis $H_{0}:X\perp\!\!\!\perp Y\mid Z$ and the alternative hypothesis $H_{1}:X\not\!\perp\!\!\!\perp Y\mid Z$, with data generated as follows:
@@ 4.1 Efficiency through Ensemble Framework
Figure 2: Comparison of Type I error (left), test power (middle), and runtime (right) for KCIT, RCIT, FastKCIT, and E-KCIT under different noise distributions.
In this experiment, we show that the ensemble framework reduces computational cost while maintaining competitive performance. We compare our ensemble-enhanced KCIT (E-KCIT) with RCIT (Strobl et al., 2019), FastKCIT (Schacht & Huang, 2025) (the only other methods known to accelerate CITs), and the original KCIT over 1000 independent trials. For all these methods, the kernel bandwidth is determined using the median heuristic for a fair comparison. Figure 2 shows results under standard Student’s $t$ (with two degrees of freedom), Cauchy, and Laplace noise, with $Z$ normally distributed, evaluating Type I error (left), test power (middle), and runtime (right).
As shown in Figure 2, E-KCIT significantly reduces computational costs while maintaining competitive test power. Notably, under the more challenging heavy-tailed noise distributions (Figures 2(a) and 2(b)), E-KCIT demonstrates more consistent performance. Although all methods generally maintain the Type I error near the nominal significance level, the ensemble framework may slightly affect this control in some scenarios. As discussed in Appendix F, this effect is independent of the test’s power. In some scenarios (as observed in subsequent experiments), can lead to a more conservative Type I error.
@@ 4.2 Ensemble Effectiveness across Conditions
Table 2: Results for $n=1200$, Standard Normal $Z$
@@
Across various simulation settings, the ensemble test consistently enhances the test power of RCIT, LPCIT, and the Fisher Z-test, while maintaining appropriate Type I error control. In contrast, the benefit for CMIknn is less pronounced. CCIT represents a special case: in our experiments, it fails to properly control the Type I error. Interestingly, applying the ensemble test significantly reduces the Type I error in this case, albeit with a minor reduction in power. We also observe that the choice of the E-CIT parameter $\alpha$ affects performance across different CIT methods and data configurations. For example, in Table 2 with $n=1200$ and $Z$ drawn from a standard normal distribution, ensemble framework with $\alpha=2$ performs better for RCIT, whereas $\alpha=1.75$ performs better for the Fisher Z-test. These observations align with our analysis in Section 3, highlighting the importance of flexibility in the E-CIT framework. We also evaluate the impact of the dimensionality of the conditioning set Z in Appendix E.3.
4.3 Real Data Experiment: Flow-Cytometry Dataset
Table 3: Performance comparison on the Flow-Cytometry dataset (results for both $\alpha$ merged).
@@ 5 Discussion
In this paper, we have introduced the Ensemble Conditional Independence Test (E-CIT), a generic, plug-and-play framework that addresses the critical computational bottleneck in constraint-based causal discovery. By employing a divide-and-aggregate strategy, E-CIT can linearizes the complexity of base CIT. Moreover, based on stable distributions, our novel p-value combination method ensures statistical properties under mild conditions. The significance of our work lies in its generality: instead of proposing another specific CIT, we present a framework that enhances the scalability and can provide consistency for a broad class of CIT methods. A detailed discussion of limitations and future directions, including Type I error behavior and the reliance on detailed properties of base CITs, is deferred to Appendix F. We believe E-CIT offers a powerful solution that achieves computational efficiency while maintaining statistical power, thereby paving the way for causal discovery in large-scale and complex scientific problems.
@@ Appendix A Closure Property of Stable Distributions
This is the general version of Proposition 1, which is one of the core properties of stable distributions. For detailed properties about stable distributions, refer to Nolan (2012; 2020).
@@ B.1 Proof of Theorem 1
The ensemble test $\mathcal{H}_{e}$ satisfies the following properties:
@@
Furthermore, since $p_{1},p_{2},\dots,p_{K}$ are derived from independent tests, it follows that:
@@
In Lemma B.1, the presence of the outer function $x^{-1}(\cdot)$ of $T_{g}$ yields an equivalent test, $x(\cdot)$ is strictly increasing. By further setting equal weights and discarding constant terms, we derive the simplified statistic
@@ B.3 Proof of Theorem 2
Consider the tangent line $l(\cdot)$ at $\mathbb{E}\left[p_{k}^{H_{1}}\right]$ for $F_{S}^{-1}$. Since $\beta=\delta=0$, $F_{S}$ is monotonically increasing and convex on $\left[0,\frac{1}{2}\right]$, and therefore $F_{S}^{-1}$ is monotonically increasing and concave on $\left[0,\frac{1}{2}\right]$. Additionally, since $\mathbb{E}\left[p_{k}^{H_{1}}\right]\leq\alpha_{e}\leq\frac{1}{2}$, it follows that $l(p)\geq F_{S}^{-1}(p)$ for $p\in\left[0,\frac{1}{2}\right]$, and $l(\cdot)$ is also monotonically increasing.
@@ Appendix C Practical Considerations of Certain Consistency Guarantees
While some CIT methods offer theoretical consistency guarantees, it is important to note that their practical performance can be compromised by the challenging nature of real-world data. We acknowledge the asymptotic guarantees, but highlight how complex data environments can limit their practical applicability. It has been shown that no single CIT can be effective in all scenarios (Shah & Peters, 2018; Kim et al., 2021), which further implies that universal consistency across all scenarios is unattainable for any single CIT.
To illustrate these practical limitations, we consider the LPCIT (Scetbon et al., 2022) as an example, focusing on how its assumptions and estimation procedures can affect its convergence speed in practice:
@@ Appendix D Certain Illustrations of the Conditions in Theorem 2
Here we consider modeling the distribution of $p_{k}^{H_{1}}$ using the Beta distribution (Heard & Rubin-Delanchy, 2018). Figure 4 illustrates an idealized case where the significance level is set to 0.05 and $p_{k}^{H_{1}}\sim\mathrm{Beta}(5,95)$, so that its expectation equals 0.05, which satisfies the first condition of Theorem 2. In this setting, the power of the test corresponds to the probability that $p_{k}^{H_{1}}<0.05$, meaning the probability of correctly rejecting the null hypothesis under the alternative. As shown by the shaded area in Figure 4, this probability is approximately 0.55. In more extreme cases, the area can approach 0.5. This indicates that the first condition of Theorem 2 can be seen as a requirement on the power of subtest under $f_{1}$, and this requirement is relatively mild.
@@ E.1 Empirical Study on the Selection of $\alpha$
Table 4: Results for $n=800$, Standard Normal $Z$
@@
Table 5: Results for $n=800$, Standard Laplace $Z$
@@
Table 6: Results for $n=1200$, Standard Laplace $Z$
@@
Table 7: Results for $n=1600$, Standard Normal $Z$
@@
Table 8: Results for $n=1600$, Standard Laplace $Z$
@@ E.3 Impact of Dimensionality of Conditioning Set $Z$
E.4 Details of Experiment on Flow-Cytometry Dataset
@@
We conduct experiments on the Flow-Cytometry dataset, a widely used benchmark for evaluating CIT methods and causal discovery algorithms (Li et al., 2023b; Sen et al., 2017; Mukherjee et al., 2020; Ng et al., 2020; Zhu et al., 2019; Mooij & Heskes, 2013). This dataset originates from the study by Sachs et al. (2005), which employed multiparameter flow cytometry to measure single-cell signaling in primary human CD4+ T cells under various interventional conditions. The study quantified 11 phosphorylated proteins and phospholipids at single-cell resolution, enabling the reconstruction of a causal graph among these biomolecules.
The dataset is available in the supplementary material of Sachs et al. (2005). The graph shown in Figure 2 of that paper represents a consensus structure derived by domain experts and is used as ground truth in our experiments. Since all data in the dataset were collected under interventions, we select the first two interventional settings that are known not to alter the underlying causal structure. The resulting data include 11 variables and 1755 samples. Extending the setup of Li et al. (2023b), we extract a comprehensive set of 50 conditionally independent and 50 conditionally dependent pairs from the consensus graph for evaluation.
Similar to Section 4.2, we compare six CIT methods and their two ensemble variants. We evaluate performance using precision, recall, and F1-score. Precision and recall are defined as $\text{TP}/(\text{TP}+\text{FP})$ and $\text{TP}/(\text{TP}+\text{FN})$, respectively. The F1-score is the harmonic mean of the two. Here, TP and TN denote correctly identified conditionally dependent and independent instances, respectively, while FP and FN denote incorrect predictions. Due to slight randomness in data partitioning within E-CIT, we repeat the process 10 times and report the average. Additionally, since RCIT (Strobl et al., 2019) inherently involves more randomness, we repeat it and its ensemble version 100 times to compute the average. Furthermore, we set the ensemble framework parameter to $K=5$, resulting in $n_{k}$ being approximately 400. As in other experiments, we evaluate E-CIT with both $\alpha=1.75$ and $\alpha=2$. As the results were identical for the two $\alpha$ values, we report them jointly in Table 3.
@@ E.5 Details of Experiment on the Application in Causal Discovery
We exclude FastKCIT (Schacht & Huang, 2025) from this comparison because its assumption that the conditioning set can be well approximated by a Gaussian mixture with $V$ components results in highly unstable performance in strongly non-Gaussian scenarios. Figure 3 reports results in terms of F1-score (left), SHD (middle), and runtime (right). SHD measures the number of edge operations required to transform the estimated graph into the ground-truth graph. To isolate the effect of conditional independence testing from that of edge orientation rules in the PC algorithm, both F1-score and SHD are computed on skeleton graphs only.

### v2

@@ 4 Experiments
In this section, we comprehensively evaluate the effectiveness of E-CIT. We first demonstrate its ability to reduce computational costs while maintaining performance (Section 4.1), followed by its broad applicability across different CIT methods (Section 4.2) and strong performance on real-world datasets (Section 4.3). Additionally, we apply E-CIT in causal discovery (Section 4.4). Other results, including the impact of subset size and the advantages of our p-value combination method for CIT, can be found in Appendices E.6 and E.7.
We conduct our synthetic experiments under the post-nonlinear model, following the setup of prior works (Zhang et al., 2011; Doran et al., 2014; Bellot and van der Schaar, 2019; Scetbon et al., 2022; Li et al., 2023a; b). Specifically, we consider the null hypothesis $H_{0}:X\perp\!\!\!\perp Y\mid Z$ and the alternative hypothesis $H_{1}:X\not\!\perp\!\!\!\perp Y\mid Z$, with data generated as follows:
@@ 4.1 Efficiency through Ensemble Framework
Figure 2: Comparison of Type I error (left; 0.05 significance level marked by solid black line), test power (middle), and runtime (right) for KCIT, RCIT, FastKCIT, and E-KCIT under different noise distributions.
In this experiment, we show that the ensemble framework reduces computational cost while maintaining competitive performance. We compare our ensemble-enhanced KCIT (E-KCIT) with RCIT (Strobl et al., 2019), FastKCIT (Schacht and Huang, 2025) (the only other methods known to accelerate CITs), and the original KCIT over 1000 independent trials. For all these methods, the kernel bandwidth is determined using the median heuristic. While advanced bandwidth optimization strategies represent an active area of research (Wang et al., 2025), we strictly adhere to the standard median heuristic as to ensure a controlled comparison and minimize potential confounding factors.
Figure 2 shows results under standard Student’s $t$ (with two degrees of freedom), Cauchy, and Laplace noise, with $Z$ normally distributed, evaluating Type I error (left), test power (middle), and runtime (right). E-KCIT significantly reduces computational costs while maintaining competitive test power. Notably, under the more challenging heavy-tailed noise distributions (Figures 2(a) and 2(b)), E-KCIT demonstrates more consistent performance. Although all methods generally maintain the Type I error near the nominal significance level, the ensemble framework may slightly affect this control in some scenarios. As discussed in Appendix F, this effect is independent of the test’s power. In some scenarios (as observed in subsequent experiments), it can lead to a more conservative Type I error.
@@ 4.2 Ensemble Effectiveness across Conditions
Table 2: Results for $n=1200$, Standard Normal $Z$. Bold (Bold italics) indicates the ensemble (original) version is statistically significantly better.
@@
Table 3: Performance comparison on the Flow-Cytometry dataset A (results for both $\alpha$ merged). Bold (Bold italics) indicates the ensemble (original) version is statistically significantly better.
@@
Across various simulation settings, the ensemble test consistently enhances the test power of RCIT, LPCIT, and the Fisher Z-test, while maintaining appropriate Type I error control. In contrast, the benefit for CMIknn is less pronounced. CCIT represents a special case: in our experiments, it fails to properly control the Type I error. Interestingly, applying the ensemble test significantly reduces the Type I error in this case, albeit with a minor reduction in power. We also observe that the choice of the E-CIT parameter $\alpha$ affects performance across different CIT methods and data configurations. For example, in Table 2 with $n=1200$ and $Z$ drawn from a standard normal distribution, the ensemble framework with $\alpha=2$ performs better for RCIT, whereas $\alpha=1.75$ performs better for the Fisher Z-test. These observations align with our analysis in Section 3, highlighting the importance of flexibility in the E-CIT framework. We also evaluate the impact of the dimensionality of the conditioning set $Z$ in Appendix E.3.
4.3 Real Data Experiment: Flow-Cytometry Dataset
@@ 5 Discussion
In this paper, we have introduced the Ensemble Conditional Independence Test (E-CIT), a general-purpose, plug-and-play framework that addresses the critical computational bottleneck in constraint-based causal discovery. By employing a divide-and-aggregate strategy, E-CIT can linearize the complexity of a base CIT. Moreover, based on stable distributions, our novel p-value combination method ensures statistical properties under mild conditions. Our theoretical and empirical findings suggest that the framework is especially effective in complex real-world scenarios. The significance of our work lies in its modularity: instead of proposing another specific CIT, we present a framework that enhances the scalability and can provide consistency for a broad class of CIT methods.
Applicability and Scope. While E-CIT provides a broadly applicable plug-and-play framework for reducing the computational burden of CITs, it is crucial to delineate its present theoretical and practical scope. First, our current analysis primarily assumes independent and identical subtests, yielding i.i.d. p-values. This condition may not be satisfied in certain scenarios (see the discussions on correlated p-values and distribution drifts in Appendix F). Second, although our p-value combination method ensures power consistency under mild conditions, its ultimate performance remains intrinsically tied to the base CIT. Consequently, while E-CIT can linearize complexity with respect to sample size, it does not resolve fundamental statistical challenges, such as the curse of dimensionality in high-dimensional conditioning sets $Z$, inherent in specific methods. Finally, the theoretical guarantees of E-CIT are conditional on the subtests being reasonably effective. As a divide-and-aggregate framework, E-CIT cannot substitute for the fundamental statistical validity of the underlying subtests.
Practical Implementation. Grounded in our theoretical analysis and empirical findings, we provide the following current guidelines for hyperparameter selection (see Appendix F for a detailed discussion). First, regarding the stability parameter $\alpha$, we recommend a general default setting of $\alpha\in\{1.75,2\}$. While our framework offers the flexibility to tune $\alpha$ for optimality relative to specific base CITs, these values have demonstrated robustness across diverse settings in our experiments. Second, regarding the subset size $n_{k}$, it should be chosen such that the base CIT exhibits reasonable asymptotic behavior and power. A practical rule of thumb is to adopt the sample size recommended in the base CIT’s original literature (e.g., $n_{k}=400$ for KCIT (Zhang et al., 2011)).
Limitations and Future Directions. We further provide a detailed discussion on the limitations and future directions of our method in Appendix F. This includes an analysis of the potential impact on Type I error control, the handling of super-uniform p-values generated by permutation tests, and promising avenues for future research such as handling correlated p-values and developing method-specific enhancements. We believe E-CIT offers a powerful solution that balances computational efficiency with statistical power, thereby paving the way for causal discovery in large-scale and complex scientific problems.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (62376243), the National Key Research and Development Program of China (2024YFE0203700), and “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2025C02037).
A preliminary version of this work was presented as the bachelor’s thesis of the first author, Zhengkang Guan. He would like to thank the faculty members of the Department of Statistics and Data Science at Xiamen University for their support and mentorship during his undergraduate studies. He is particularly grateful to Prof. Jingyuan Liu for her insightful discussions on general statistical methodologies, as well as her invaluable guidance regarding overall academic and career development.
@@ Appendix A Closure Property of Stable Distributions
This is the general version of Proposition 1, which is one of the core properties of stable distributions. For detailed properties of stable distributions, refer to Nolan (2012; 2020).
@@ B.1 Proof of Theorem 1
The ensemble test $\mathcal{H}_{e}$ (for exact subtest p-values) satisfies the following properties:
@@
Furthermore, since $p_{1},p_{2},\dots,p_{K}$ are derived from independent tests (and are thus i.i.d.), it follows that:
@@
In Lemma B.1, the presence of the outer function $x^{-1}(\cdot)$ of $T_{g}$ yields an equivalent test, since $x(\cdot)$ is strictly increasing. By further setting equal weights and discarding constant terms, we derive the simplified statistic
@@ B.3 Proof of Theorem 2
Consider the tangent line $l(\cdot)$ of $F_{S}^{-1}$ at $\mathbb{E}\left[p_{k}^{H_{1}}\right]$. Since $\beta=\delta=0$, $F_{S}$ is monotonically increasing and convex on $\left[0,\frac{1}{2}\right]$, and therefore $F_{S}^{-1}$ is monotonically increasing and concave on $\left[0,\frac{1}{2}\right]$. Additionally, since $\mathbb{E}\left[p_{k}^{H_{1}}\right]\leq\alpha_{e}\leq\frac{1}{2}$, it follows that $l(p)\geq F_{S}^{-1}(p)$ for $p\in\left[0,\frac{1}{2}\right]$, and $l(\cdot)$ is also monotonically increasing.
@@ Appendix C Practical Considerations of Certain Consistency Guarantees
While some CIT methods offer theoretical consistency guarantees, it is important to note that their practical performance can be compromised by the challenging nature of real-world data. We acknowledge the asymptotic guarantees, but highlight how complex data environments can limit their practical applicability. It has been shown that no single CIT can be effective in all scenarios (Shah and Peters, 2018; Kim et al., 2021), which further implies that universal consistency across all scenarios is unattainable for any single CIT.
To illustrate these practical limitations, we consider LPCIT (Scetbon et al., 2022) as an example, focusing on how its assumptions and estimation procedures can affect its convergence speed in practice:
@@ Appendix D Certain Illustrations of the Conditions in Theorem 2
Here we consider modeling the distribution of $p_{k}^{H_{1}}$ using the Beta distribution (Heard and Rubin-Delanchy, 2018). Figure 4 illustrates an idealized case where the significance level is set to 0.05 and $p_{k}^{H_{1}}\sim\mathrm{Beta}(5,95)$, so that its expectation equals 0.05, which satisfies the first condition of Theorem 2. In this setting, the power of the test corresponds to the probability that $p_{k}^{H_{1}}<0.05$, meaning the probability of correctly rejecting the null hypothesis under the alternative. As shown by the shaded area in Figure 4, this probability is approximately 0.55. In more extreme cases, the area can approach 0.5. This indicates that the first condition of Theorem 2 can be seen as a requirement on the power of the subtest under $f_{1}$, and this requirement is relatively mild.
@@ E.1 Empirical Study on the Selection of $\alpha$
Table 4: Results for $n=800$, Standard Normal $Z$. Bold (Bold italics) indicates the ensemble (original) version is statistically significantly better.
@@
Table 5: Results for $n=800$, Standard Laplace $Z$. Bold (Bold italics) indicates the ensemble (original) version is statistically significantly better.
@@
Table 6: Results for $n=1200$, Standard Laplace $Z$. Bold (Bold italics) indicates the ensemble (original) version is statistically significantly better.
@@
Table 7: Results for $n=1600$, Standard Normal $Z$. Bold (Bold italics) indicates the ensemble (original) version is statistically significantly better.
@@
Table 8: Results for $n=1600$, Standard Laplace $Z$. Bold (Bold italics) indicates the ensemble (original) version is statistically significantly better.
@@ E.3 Impact of Dimensionality of Conditioning Set $Z$
E.4 Details of Experiment on Real-World Flow-Cytometry Data
@@
We conduct experiments on the Flow-Cytometry dataset, a widely used benchmark for evaluating CIT methods and causal discovery algorithms (Li et al., 2023b; Sen et al., 2017; Mukherjee et al., 2020; Ng et al., 2020; Zhu et al., 2019; Mooij and Heskes, 2013). This dataset originates from the seminal study by Sachs et al. (2005), which employed multiparameter flow cytometry to measure single-cell signaling in primary human CD4+ T cells. The dataset is available in the supplementary material of Sachs et al. (2005). (It is important to clarify that this dataset consists of real-world biological data, not computer simulations. The term “stimulations” in the original study refers to biological reagents used to perturb the cells, not synthetic data generation.)
The dataset comprises measurements of 11 phosphorylated proteins and phospholipids. The consensus causal graph, derived by domain experts and validated in the original study, serves as the ground truth. Although the data were collected under biological stimulations, we follow the standard protocol in CIT literature (Li et al., 2023b; Sen et al., 2017; Mukherjee et al., 2020; Ng et al., 2020; Zhu et al., 2019; Mooij and Heskes, 2013) by treating specific subsets as observational data. This is justified because the stimulations act on unobserved exogenous root nodes, preserving the downstream causal structure among the observed proteins.
To provide a comprehensive evaluation and address different sample size regimes, we utilize two configurations of this dataset:
  • Dataset A ($n=1755$): Following the setup of Li et al. (2023b), we combine the samples from the cd3cd28 and cd3cd28icam2 experimental conditions. This results in a total of 1755 samples. This setting tests the methods’ performance on a moderate sample size.
  • Dataset B ($n=853$): To align with benchmarks such as CCMI (Mukherjee et al., 2020) and evaluate performance on smaller sample sizes, we explicitly evaluate on the cd3cd28 alone, which contains 853 samples.
From the consensus graph, we extract a comprehensive set of 50 conditionally independent and 50 conditionally dependent pairs for evaluation. Similar to Section 4.2, we compare six CIT methods and their two ensemble variants. We evaluate performance using precision, recall, and F1-score. Precision and recall are defined as $\text{TP}/(\text{TP}+\text{FP})$ and $\text{TP}/(\text{TP}+\text{FN})$, respectively. The F1-score is the harmonic mean of the two. Here, TP and TN denote correctly identified conditionally dependent and independent instances, respectively, while FP and FN denote incorrect predictions.
Consistent with our hyperparameter guideline to maintain a subset size $n_{k}\approx 400$, we set $K=5$ for Dataset A ($n_{k}\approx 351$) and $K=2$ for Dataset B ($n_{k}\approx 426$). We evaluate E-CIT with both $\alpha=1.75$ and $\alpha=2$. As the results were stable across these values, we report them jointly. Due to the randomness in data partitioning, E-CIT results are averaged over 10 runs. RCIT (Strobl et al., 2019) and its ensemble version are averaged over 100 runs to account for their inherent randomness. The results for the two datasets are presented in Tables 3 and 9.
Table 9: Performance comparison on the Flow-Cytometry dataset B (results for both $\alpha$ merged). Bold (Bold italics) indicates the ensemble (original) version is statistically significantly better.
Method Precision Recall F1-score
KCIT Orig. 0.880 0.667 0.759
Ensemble 0.862 0.660 0.747
RCIT Orig. 0.888 0.654 0.753
Ensemble 0.872 0.665 0.754
LPCIT Orig. 0.912 0.660 0.766
Ensemble 0.924 0.673 0.778
CMIknn Orig. 0.920 0.667 0.773
Ensemble 0.904 0.667 0.767
CCIT Orig. 0.560 0.636 0.596
Ensemble 0.594 0.658 0.623
FisherZ Orig. 0.940 0.671 0.783
Ensemble 0.942 0.672 0.784
@@ E.5 Details of Experiment on the Application in Causal Discovery
We exclude FastKCIT (Schacht and Huang, 2025) from this comparison because its assumption that the conditioning set can be well approximated by a Gaussian mixture with $V$ components results in highly unstable performance in strongly non-Gaussian scenarios. Figure 3 reports results in terms of F1-score (left), SHD (middle), and runtime (right). SHD measures the number of edge operations required to transform the estimated graph into the ground-truth graph. To isolate the effect of conditional independence testing from that of edge orientation rules in the PC algorithm, both F1-score and SHD are computed on skeleton graphs only.